Determine the density of N2 gas at STP. Molar mass N2 is 28 g/mol.
1
Expert's answer
2018-10-24T08:52:09-0400

Answer:

PV=nRT, n=m/Mr, PV=mRT/Mr, therefore P=mRT/MrV

m/V = D (density)

P=DRT/Mr D= MrP/RT

D=28 x 1 / 0.082 x 22.4 = 15.244

D= 15.244 (N 2 )